Current and Quick Ratios The Nelson Company has $1,260,000 in current assets and $450,000 in current liabilities. Its initial inventory level is $225,000, and it will raise funds as additional notes payable and use them to increase inventory.

1.How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 2.1? Round your answer to the nearest cent. $

2.What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Round your answer to two decimal places.
Answer #1

Any Current assas = Current liability - 1260000 450000 Current Retio = Current assets Current liability 1260000 450000 = 2.82.1 (450000+)) = 1200000 +X 945000 + 2) = 1200000 +X +315000 = 71.12 2 = 286363.636 Notes payable = 286363. 64 Notes payable